OpenStudy (anonymous):

Can someone solve the inequality w/ steps ? -4w+12≥10 or 5w-14>4

OpenStudy (anonymous):

Solving inequalities is not that different than solving equations. For \[-4w + 12 \ge 10 \] you would subtract 12 from both sides for \[-4w \ge -2\] which would be \[w \le 1/2\] because when you divide an inequality by a negative number you must switch the direction of the sign.